The average time taken for the nearest Covid-19 ambulance team to convey a patient showing symptoms to the nearest designated hospital follows the normal distribution about a mean of 60 seconds and a standard deviation of 8 seconds. If 200 patients were selected at random, how many of them could be taken to the hospital in i) Less than 50 seconds (3 Marks) ii) More than 64 seconds (
1
Expert's answer
2020-06-07T13:26:33-0400

"ii) Given \; that, μ=60, σ=8, n=200, then,\\
 P(x<50)=P(Z<\frac{50-60}{8})\\
=P(Z<-1.25)=0.5-P(0<Z<1.25)\\
=0.5-0.3944=0.1056\\
 number\;of\;patients=200\times0.1056=21.12\approx22\\
ii) P(x>64)=P(Z>\frac{64-60}{8})\\
=P(Z>0.5)=0.5-P(0<Z<0.5)\\
=0.5-0.1915=0.3085\\
 number\;of\;patients=200\times0.3085=61.7\approx62\\"